# Coded Downlink Massive Random Access and a Finite de Finetti Theorem

### v1

@@ I-C Main Contributions
I-C1 Coding for Downlink Massive Random Access
For transmitting an exchangeable source $\mathbf{X}=(X_{1},\cdots,X_{k})\sim p(\mathbf{x})$ to a random subset of $k$ active users in a downlink massive random access scenario with $n$ total users, we show that a common message rate of
$\min_{q\in\mathcal{Q}}H(\mathbf{X})+D(p\|q)+\log(H(\mathbf{X})+D(p\|q)+1)+1$ (4)
bits is achievable, where $\mathcal{Q}$ is the family of i.i.d. mixture distributions on $\mathcal{X}^{k}$ and $D(\cdot\|\cdot)$ is the Kullback-Leibler (KL) divergence.
In the case that the source is i.i.d. or i.i.d. mixture, we can set $q=p$ to eliminate the $D(p\|q)$ term. Specifically, for i.i.d. sources, an achievable rate is $kH(X_{1})+\log(kH(X_{1})+1)+1$, thus the overhead is at most $O(\log(k))$ bits. The overhead can be reduced to $O(1)$ if the source is uniform over its support.
@@ I-C1 Coding for Downlink Massive Random Access
$\min_{q\in\mathcal{Q}}D(p\|q)\leq\min\{k\log(e),|\mathcal{X}|\log(k+1)\}.$ (5)
thus achieving an overhead of at most $O(k)$ bits in general and $O(\log(k))$ bits if the source has a finite alphabet. In all cases, the achievable common message rate is close to $H(\mathbf{X})$ bits with an overhead that does not depend on $n$.
I-C2 Extendable Sources and Finite de Finetti Theorem
The upper bound (5) can be improved if $\mathbf{X}\sim p(\mathbf{x})$ is $d$-extendable. In this case, we show that
$\min_{q\in\mathcal{Q}}D(p\|q)\leq\min\left\{(|\mathcal{X}|-1)\log\left(\frac{d-1}{d-k}\right),\right.\\ \left.-\log\left(1-\frac{k(k-1)}{2d}\right)\right\}.$ (6)
This result is a statement of the finite de Finetti theorem in terms of relative entropy. In particular, for any fixed $k$, the relative entropy between the distribution $p(\mathbf{x})$ and the nearest i.i.d. mixture distribution goes to 00 as $d\to\infty$. In fact, the scaling in (6) matches the optimal scaling of (2), because the upper bound goes to 00 as $k,d\to\infty$, as long as $\frac{k^{2}}{d}\to 0$, or for the finite alphabet case, $\frac{k}{d}\to 0$.
I-D Paper Organization and Notation
@@
Throughout this paper, we use lowercase letters to denote scalars, lowercase boldface letters to denote vectors, capital letters to denote random variables, boldface capital letters to denote random vectors, and calligraphic letters, i.e. $\mathcal{S}$, to denote sets, $|\mathcal{S}|$ to denote their cardinality, and $\mathbbm{1}_{\mathcal{S}}$ as the indicator function. We let $\log(\cdot)$ denote the base 2 logarithm and $\ln(\cdot)$ denote the natural logarithm. All information measures are expressed in bits, including entropy $H(\cdot)$ and the relative entropy or KL divergence $D(\cdot\|\cdot)$. We use the short-hands notations $[n]=\{1,\dots,n\}$ and $a^{\underline{b}}=a(a-1)\cdots(a-b+1)$.
@@ II-A Problem Formulation
Consider a massive random access setting in which a random subset of $k$ users becomes active among a total number of $n$ users. The value of $n$ is fixed and known. For now, $k$ is also assumed to be fixed and known. The identities of the active users are known to the BS. However, each user only knows whether itself is active, but it does not know which other users are active. We are interested in the setting in which the BS wishes to communicate a source to each of the $k$ active users simultaneously using a common broadcast message. This common message is assumed to be received by all users without error.
In this paper, we consider the class of source distributions that are exchangeable. Let $\mathbf{X}=(X_{1},\cdots,X_{k})$ be a sequence of exchangeable random variables taking values in a discrete alphabet $\mathcal{X}$. Let $p(\mathbf{x})=p(x_{1},\cdots,x_{k})$ be their joint distribution. Each user $i$ is interested in learning $X_{i}$ based on the common message (but not interested in any of other $X_{i}$’s).
@@
    The $k$ active users need to be assigned into $b$ slots, where $b\geq k$, in a non-colliding fashion, i.e., subject to the constraint that no two users can be assigned to the same slot. This is known as scheduling. Fixing a particular schedule, the BS wishes to communicate to each user which slot it is being assigned to.
@@
    The $k$ active users need to be categorized into $c$ categories, but each category must have a fixed number of $k_{\ell}$ users, $\forall\ell\in[c]$, $\sum_{\ell=1}^{c}k_{\ell}=k$. The BS wishes to communicate the category labels $X_{i}\in[c]$ to each user. For the case of two categories ($c=2$), this is known as the acknowledgement problem.
@@
    A fixed number of resources $r$ need to be distributed amongst the $k$ active users, i.e., $\sum_{i=1}^{k}X_{i}=r$, where $X_{i}\geq 0$ is an integer. The BS wishes to communicate to each user how many units of resources it has been assigned. This is known as the resource allocation problem.
@@
Here, $a_{i}\in[n]$ is the index of the $i$th active user. While the source $\mathbf{X}$ describes the contents of the messages, the activity pattern $\mathbf{A}$ indicates which users should receive which message, i.e., we want each user $a_{i}$ to receive $x_{i}$ $\forall i\in[k]$. Together, they form a message-activity tuple $(\mathbf{X},\mathbf{A})$. Throughout this paper, we assume that $\mathbf{X}$ and $\mathbf{A}$ are independent. Notationally, we use $(\mathbf{x},\mathbf{a})$ to represent a realization of $(\mathbf{X},\mathbf{A})$.
The problem of communicating the sources $\mathbf{X}$ to the active users in $\mathbf{A}$ can now be thought of as a one-shot source coding problem consisting of a single encoder and multiple decoders. In this paper, we pursue the following two-stage strategy for encoding. In the first stage, the BS uses encoder $f$ to map a message-activity pair $(\mathbf{x},\mathbf{a})$ to a natural number, i.e.,
$f:\mathcal{X}^{k}\times\mathcal{A}^{(n,k)}\rightarrow\mathbb{N}.$ (8)
In the second stage, entropy coding is used to compress $f(\mathbf{x},\mathbf{a})$ into a variable-length codeword which is then broadcast to all active users.
On the decoding side, assuming error-free broadcasting of the common message, each active user $a_{i}$ first recovers $f(\mathbf{x},\mathbf{a})$. It then uses its own decoder $d_{a_{i}}:\mathbb{N}\to\mathcal{X}$ to recover the intended message. We require lossless recovery:
@@
i.e. each user receives its intended source without error.
The rate of the encoding and decoding scheme is defined as $R=H(f(\mathbf{X},\mathbf{A}))$, where $H(\cdot)$ is the entropy function, justified by the fact that entropy coding can be used to achieve the rate $R$ to within one bit.
The optimal encoding and decoding scheme is defined to be the encoder and decoders that minimize $H(f(\mathbf{X},\mathbf{A}))$, while satisfying the condition (9). The optimal common message rate is defined as
$R^{*}\triangleq H\left(f^{*}(\mathbf{X},\mathbf{A})\right),$ (10)
@@ II-B Codebook Construction
We propose the following encoding and decoding scheme inspired by previous work [6] [1] for coded downlink massive random access. This encoding and decoding scheme utilizes a shared codebook between the BS and all the users
$\mathbf{m}=(\mathbf{c}^{(1)},\mathbf{c}^{(2)},\cdots)$ (11)
which consists of an infinite sequence of length-$n$ vectors $\mathbf{c}^{(t)}\in\mathcal{X}^{n}$. We assign a unique entry location in the codewords to each of the $n$ users. For a particular message-activity tuple $(\mathbf{x},\mathbf{a})$, the BS encodes $(\mathbf{x},\mathbf{a})$ by finding a codeword $\mathbf{c}^{(t)}$ such that every active user has the correct message in their designated entry in $\mathbf{c}^{(t)}$, i.e.,
@@
Since the codebook is of infinite size and $\mathcal{X}$ is a discrete alphabet, such a codeword always exists. The idea is to encode $(\mathbf{x},\mathbf{a})$ as the index of the first such codeword in the codebook. Mathematically, the proposed encoder can be written as
$f_{\mathbf{m}}(\mathbf{x},\mathbf{a})=$ $\min$ $t$
@@
The decoders for each user $u\in[n]$ are defined as:
@@
It remains to discuss how to generate the codewords in $\mathbf{m}$. Similar to [1], we use a random codebook construction in which the codewords are generated independently according to a mixture of i.i.d. distributions. Specifically, fix a distribution $r(\theta)$ and a conditional distribution $q(\mathbf{x}|\theta)$. Each length-$n$ codeword $\mathbf{c}^{(t)}=[x_{1},\cdots,x_{n}]^{\top}$ is generated according to
@@
Operationally, each codeword can be thought of as being generated in a two-step process of first choosing a $\theta$ according to $r(\theta)$, then generating the entries of the codeword in an i.i.d. fashion according to $q(\mathbf{x}|\theta)$. Each subsequent codeword is generated independently in the same way. Once the entire codebook is generated, it is shared with the BS and all users.
@@
This means that the distribution of any $k$ distinct entries of the codewords are identical, which is a desired property for encoding exchangeable sources. Further, this distribution can be specifically designed according to any i.i.d. mixture.
This paper focuses attention to the encoding and decoding functions and random codebooks generated in this way. Our main result is that for any exchangeable source distribution $p(x_{1},\cdots,x_{k})$, we can choose an appropriate i.i.d. mixture distribution $q(\mathbf{x}|\theta)r(\theta)$, such that among an ensemble of codebooks randomly generated according to (15), there exists a codebook $\mathbf{m}^{*}$ that has an encoder output entropy $H(f_{\mathbf{m}^{*}}(\mathbf{X},\mathbf{A}))$ upper bounded by $H(X_{1},\cdots,X_{k})$ plus a small overhead term.
II-C Achievable Rates
In this section, we provide upper bounds on the optimal rate $R^{*}$ of the downlink message for massive random access by analyzing the entropy of the output of the encoder when
  • The source ${\mathbf{X}}=(X_{1},\cdots,X_{k})$ is distributed according to an exchangeable source distribution $p(x_{1},\cdots,x_{k})$;
  • The codebook is generated according to i.i.d. mixture distribution $q(\mathbf{x}|\theta)r(\theta)$.
@@ II-C Achievable Rates
Consider a massive random access scenario with a total of $n$ users and a random subset of $k$ active users. Let sources $\mathbf{X}=(X_{1},\cdots,X_{k})$ take values in a discrete set $\mathcal{X}^{k}$ with exchangeable distribution $p(\mathbf{x})$. The minimum common message rate $R^{*}$ is bounded above as
$R^{*}\leq\min_{q\in\mathcal{Q}}H(\mathbf{X})+D(p\|q)+\log(H(\mathbf{X})+D(p\|q)+1)+1\$ (17)
where $\mathcal{Q}$ is the family of all i.i.d. mixture distributions on $\mathcal{X}^{k}$. defined as in (15). Further, the minimum common message rate $R^{*}$ is also bounded above as:
$R^{*}\leq\min_{q\in\mathcal{Q}}H(\mathbf{X})+D(p\|q)+\log\left(\log\left(\frac{q_{\text{max}}}{q_{\text{min}}}\right)+1\right)+3,$ (18)
@@
$q_{\text{max}}=\max_{\mathbf{x}\in\mathcal{X}^{k}:p(\mathbf{x})>0}q(\mathbf{x});\quad q_{\text{min}}=\min_{\mathbf{x}\in\mathcal{X}^{k}:p(\mathbf{x})>0}q(\mathbf{x}).$ (19)
@@
The key insight of Theorem 1 is the $D(p\|q)$ term, which corresponds to the cost of using a codebook constructed according to $q({\mathbf{x}})$ to compress a source with distribution $p({\mathbf{x}})$. This is the same $D(p\|q)$ cost as in classical source coding. For i.i.d. mixture sources, the $D(p\|q)$ cost can be eliminated.
@@
Under the setting of Theorem 1, if the source $\mathbf{X}$ is i.i.d. or a mixture of i.i.d.  then the minimum common message rate is bounded above by
$R^{*}\leq H(\mathbf{X})+\log(H(\mathbf{X})+1)+1.$ (20)
Further, for i.i.d. sources with distribution $p(x)$, we have
$R^{*}\leq kH(p)+\min\bigg{\{}\log(kH(p)+1)+1,\\ \left.\log\left(k\log\left(\frac{p_{\text{max}}}{p_{\text{min}}}\right)+1\right)+3\right\},$ (21)
@@
$p_{\text{max}}=\max_{{x}\in\mathcal{X}:p({x})>0}p({x});\quad p_{\text{min}}=\min_{{x}\in\mathcal{X}:p({x})>0}p({x}).$ (22)
@@
Setting $q(\mathbf{x})=p(\mathbf{x})$ is a natural choice for i.i.d. or i.i.d. mixture sources, because it maximizes the probability of matching codewords with the source. This result shows that for i.i.d. and i.i.d. mixture sources, the proposed code construction and the proposed encoding and decoding scheme can achieve the source entropy to within at most an $O(\log(k))$ overhead. The overhead can be further reduced if the distribution is close to uniform. In particular, for an i.i.d. uniform distribution with $p_{\text{max}}=p_{\text{min}}$, the entropy can be achieved to within $O(1)$ bits.
An intuitive (but somewhat inaccurate) argument of why for i.i.d. sources, a matching i.i.d. codebook can approach the source entropy is the following. When the codebook distribution and the source distribution $p(\mathbf{x})$ match, for every realization of the source $\mathbf{x}$, a randomly generated codeword would match $\mathbf{x}$ with a probability $p(\mathbf{x})$. Thus, the index of the first matching codeword follows a geometric distribution with parameter $p(\mathbf{x})$. The entropy of such a geometric distribution is essentially $-\log(p(\mathbf{x}))$ plus a constant. Then, when averaged over all realizations of $\mathbf{x}$, we can achieve a common message entropy of $\mathbb{E}[-\log(p(\mathbf{X}))]$, which is just the source entropy $H(\mathbf{X})$, plus a constant.
To make the above argument rigorous, we recognize that the actual distribution of the first matching index for a fixed codebook, across all source realizations, is a mixture of geometric distributions, for which the computation of entropy is quite nontrivial. Appendix A uses a bounding technique over randomly generated codebooks to show the existence of one codebook whose output has the desired entropy, up to an additive logarithmic term.

### v4

@@ I-E Main Contributions
I-E1 Coding for Downlink Massive Random Access
For transmitting exchangeable sources $\mathbf{X}=(X_{1},\cdots,X_{k})\sim p(\mathbf{x})$ to a random subset of $k$ active users in a downlink massive random access scenario with $n$ total users, we show that a common message length of
$\min_{q\in\mathcal{Q}}H(\mathbf{X})+D(p\|q)+\log(H(\mathbf{X})+D(p\|q)+1)+2$ (10)
bits is achievable, where $\mathcal{Q}$ is the family of i.i.d. mixture distributions on $\mathcal{X}^{k}$.
In the case that the sources are i.i.d. or i.i.d. mixture, we can set $q(\mathbf{x})=p(\mathbf{x})$ to eliminate the $D(p\|q)$ term. Specifically, for i.i.d. sources, an achievable common message length is $kH(X_{1})+\log(kH(X_{1})+1)+2$, thus the overhead is at most $O(\log(k))$ bits. The overhead can be reduced to $O(1)$ if the sources are uniform over the alphabet in its support.
@@ I-E1 Coding for Downlink Massive Random Access
$\min_{q\in\mathcal{Q}}D(p\|q)\leq\min\{k\log(e),|\mathcal{X}|\log(k+1)\},$ (11)
thus achieving an overhead of at most $O(k)$ bits in general and $O(\log(k))$ bits if the source has a finite alphabet.
In all cases, the achievable common message length is close to $H(\mathbf{X})$ bits with an overhead that does not depend on $n$.
I-E2 Extendable Sources and Finite de Finetti Theorem
The upper bound (11) can be improved if the sources $\mathbf{X}\sim p(\mathbf{x})$ are $d$-extendable. In this case, we show that for $d>k$
$\min_{q\in\mathcal{Q}}D(p\|q)\leq\min\left\{\log\left(\frac{d^{k}}{d^{\underline{k}}}\right),(|\mathcal{X}|-1)\log\left(\frac{d-1}{d-k}\right)\right\},$ (12)
where $d^{\underline{k}}=d(d-1)\cdots(d-k+1)$. This is a statement of the finite de Finetti theorem in terms of KL divergence.
This bound is independent of the distribution of the sources. Further, by using the fact $\log\left(\frac{d^{k}}{d^{\underline{k}}}\right)\leq-\log\left(1-\frac{k(k-1)}{2d}\right)$ for $d>\frac{1}{2}k(k-1)$, it can be shown that the KL divergence between the distribution $p(\mathbf{x})$ and the nearest i.i.d. mixture distribution goes to 00 as long as $\frac{k^{2}}{d}\to 0$ for general alphabet case and $\frac{k}{d}\to 0$ in the finite alphabet case. These scaling results match the optimal scalings of (7).
I-F Paper Organization and Notation
@@
Throughout this paper, we use lowercase letters to denote scalars, lowercase boldface letters to denote vectors, capital letters to denote random variables, boldface capital letters to denote random vectors, and calligraphic letters, e.g. $\mathcal{S}$, to denote sets, $|\mathcal{S}|$ to denote their cardinality. We let $\log(\cdot)$ denote the base 2 logarithm and $\ln(\cdot)$ denote the natural logarithm. All information measures are expressed in bits, including entropy $H(\cdot)$ and the KL divergence $D(\cdot\|\cdot)$. We use $\mathbb{N}$ to denote the set of natural numbers and $\mathbbm{1}(\cdot)$ as the indicator function. We use Ber$(p)$ to denote Bernoulli random variable with parameter $p$. We use the shorthand notations $[n]=\{1,\dots,n\}$ and $a^{\underline{b}}=a(a-1)\cdots(a-b+1)$.
@@ II-A Problem Formulation
Consider a massive random access setting in which a random subset of $k$ users become active among a total number of $n$ users. The value of $n$ is fixed and known. For now, $k$ is also assumed to be fixed and known. The identities of the active users are known to the BS. However, each user only knows whether itself is active, but does not know which other users are active. We are interested in the setting where the BS wishes to communicate a source to each of the $k$ active users simultaneously using a common message in the downlink. This common message is assumed to be received by all users without error.
In this paper, we consider the class of sources that are exchangeable. Let $\mathbf{X}=(X_{1},\cdots,X_{k})$ be a sequence of exchangeable random variables taking values from a finite or countable alphabet $\mathcal{X}$. Let $p(\mathbf{x})=p(x_{1},\cdots,x_{k})$ be their joint distribution. Each user $i$ is interested in learning $X_{i}$ based on the common message (but not interested in any of other $X_{i}$’s).
@@
    The $k$ active users need to be assigned into $b$ slots, where $b\geq k$, in a non-colliding fashion, i.e., no two users can be assigned to the same slot. This is known as scheduling. Fixing a particular schedule, the BS wishes to communicate to each user which slot it is assigned to.
@@
    The $k$ active users need to be categorized into $c$ categories where each category has a fixed number of $k_{\ell}$ users and $\sum_{\ell=1}^{c}k_{\ell}=k$. The BS wishes to communicate the category labels $X_{i}\in[c]$ to each user. For the case of two categories ($c=2$), this is known as the acknowledgment problem.
@@
    A fixed number of resources $r$ need to be distributed among the $k$ active users, i.e., $\sum_{i=1}^{k}X_{i}=r$, where $X_{i}\geq 0$ is an integer. The BS wishes to communicate to each user how many units of resources it has been assigned. This is known as the resource allocation problem.
@@
Here, $a_{i}\in[n]$ is the index of the $i$th active user. While $\mathbf{X}$ describes the source contents, the activity pattern $\mathbf{A}$ indicates which users should receive which source, i.e., we want each user $a_{i}$ to receive $x_{i}$ $\forall i\in[k]$. Together, they form a source-activity pair $(\mathbf{X},\mathbf{A})$. Throughout this paper, we assume that $\mathbf{X}$ and $\mathbf{A}$ are independent. Notationally, we use $(\mathbf{x},\mathbf{a})$ to represent a realization of $(\mathbf{X},\mathbf{A})$.
The problem of communicating the sources $\mathbf{X}$ to the active users in $\mathbf{A}$ can now be thought of as a one-shot source coding problem consisting of a single encoder and multiple decoders. The BS uses an encoder $f$ to map a source-activity pair $(\mathbf{x},\mathbf{a})$ to a binary string, i.e.,
$f:\mathcal{X}^{k}\times\mathcal{A}^{(n,k)}\rightarrow\{0,1\}^{*}.$ (14)
It is assumed that the binary string is sent to all the active users in an error-free fashion. Each active user then uses its decoder $d_{a_{i}}:\{0,1\}^{*}\to\mathcal{X}$ to recover its intended source based on the binary string. We require lossless recovery of the sources:
@@
i.e. each active user must receive its intended source without error. Note that users are not required to learn anything pertaining to the other active users, whether that be their identities or the intended sources for the other users.
To ensure unique decodability, we require the set of all possible output binary strings of the encoder to be a prefix-free code. The optimal encoder and decoders $(f^{*}(\mathbf{x},\mathbf{a}),d_{a_{1}}^{*},\cdots d_{a_{k}}^{*})$, $(\mathbf{x},\mathbf{a})\in\mathcal{X}^{k}\times\mathcal{A}^{(n,k)}$ are defined to be the encoder and decoders that minimize $\mathbb{E}[\mathtt{len}(f(\mathbf{X},\mathbf{A}))]$ with $\mathtt{len}(\cdot)$ denoting the length of a string, while satisfying the condition (15). The optimal common message length is
$R^{*}\triangleq\mathbb{E}\left[\mathtt{len}\left(f^{*}(\mathbf{X},\mathbf{A})\right)\right],$ (16)
@@
Remark 1
This paper focuses on the setting in which the downlink channel is noiseless and has the same capacity to all users. This turns the design of optimal common message encoder and decoders into a source coding problem. If instead, the downlink channel is noisy, then either appropriate channel coding can be used to turn the downlink channel into an error-free channel, or if the problem is analyzed from the point of view of transmitting correlated sources over a noisy broadcast channel (e.g., [15]), it would lead to a highly complex problem setting, which we will not treat in this paper.
Remark 2
The definitions of the encoder (14) and the expected codeword length (16) assume that the common message can have variable length. Although in many communications context, fixed-length codes may be more desirable at a protocol level, the optimal fixed-length codes are considerably more difficult to construct for this downlink massive random access problem. The existence of fixed-length codes has been shown for specific problem instances such as collision-free scheduling [6], categorization [1], and user acknowledgment [16], but the proofs often involve the use of probabilistic arguments and are non-constructive [17]. Further, if the source alphabet is infinite, fixed-length codes cannot exist. For these reasons, the rest of this paper focuses on code construction for variable-length common messages.
@@ II-B Codebook Construction
This paper proposes an encoding and decoding strategy involving two stages. The BS first uses a function $g$ to map a source-activity pair to a positive index, i.e.,
$g:\mathcal{X}^{k}\times\mathcal{A}^{(n,k)}\rightarrow\mathbb{N},$ (17)
then uses an optimal prefix-free code to compress the output of $g(\mathbf{x},\mathbf{a})$ into a variable-length binary string $f(\mathbf{x},\mathbf{a})$, which is subsequently broadcast to all active users. On the decoding side, each active user $a_{i}$ first decodes $g(\mathbf{x},\mathbf{a})$ based on the received binary string, then recovers their respective sources.
Specifically, the proposed encoding and decoding scheme utilizes a shared codebook between the BS and all the users
$\mathbf{m}=(\mathbf{c}^{(1)},\mathbf{c}^{(2)},\cdots),$ (18)
which consists of a potentially infinite sequence of length-$n$ vectors $\mathbf{c}^{(t)}\in\mathcal{X}^{n}$. We assign a unique entry location in the codewords to each of the $n$ users. For a particular source-activity tuple $(\mathbf{x},\mathbf{a})$, the BS encodes $(\mathbf{x},\mathbf{a})$ by finding a codeword $\mathbf{c}^{(t)}$ such that every active user has the correct source content in their designated entry in $\mathbf{c}^{(t)}$, i.e.,
@@
We construct the codebook $\mathbf{m}$ so that such a matching codeword always exists. If there are multiple matching codewords, we define the encoder output to be the index of the first matching codeword in the codebook.
Denoting $g_{\mathbf{m}}(\mathbf{x},\mathbf{a})$ as the encoder output for the source-activity pair $(\mathbf{x},\mathbf{a})$ using the codebook $\mathbf{m}$, we have
$g_{\mathbf{m}}(\mathbf{x},\mathbf{a})=$ $\min$ $t$
@@
This output is further compressed using a prefix-free code. Since any random variable can be compressed using an optimal prefix-free code into a binary string with an expected length within one bit of its entropy, the common message length of the proposed encoding scheme is bounded by
$R<H(g_{\mathbf{m}}(\mathbf{X},\mathbf{A}))+1.$ (21)
At the decoder, each user decompresses the prefix-free code, thereby recovering $g_{\mathbf{m}}(\mathbf{x},\mathbf{a})$ or equivalently the minimum $t$. The decoding process for each user $u\in[n]$ is simply
@@
It remains to discuss how to generate the codewords in $\mathbf{m}$. We use a random codebook construction in which the codewords are generated independently according to a mixture of i.i.d. distributions. Specifically, fix a distribution $r(\theta)$ and a conditional distribution $q(\mathbf{x}|\theta)$. Each length-$n$ codeword $\mathbf{c}^{(t)}=[x_{1},\cdots,x_{n}]^{\top}$ is generated according to
@@
Operationally, each codeword can be thought of as being generated in a two-step process of first choosing a $\theta$ according to $r(\theta)$, then generating the entries of the codeword in an i.i.d. fashion according to $q(\mathbf{x}|\theta)$. This makes each entry of the codeword conditionally i.i.d. Each subsequent codeword is generated independently in the same way and the entire codebook is shared between the BS and all users.
@@
This means that the distributions of any $k$ distinct entries of the codewords are identical, which is a desired property for encoding exchangeable sources. This distribution can be specifically designed according to any i.i.d. mixture.
This paper focuses attention to the encoding and decoding functions and random codebooks generated in this way. The main result of this paper is that for any exchangeable distribution $p(x_{1},\cdots,x_{k})$, we can choose an appropriate i.i.d. mixture distribution $q(\mathbf{x}|\theta)r(\theta)$, such that among an ensemble of codebooks randomly generated according to (23), there exists a codebook $\mathbf{m}^{*}$ that has a corresponding $H(g_{\mathbf{m}^{*}}(\mathbf{X},\mathbf{A}))$ upper bounded by $H(X_{1},\cdots,X_{k})$ plus a small overhead.
Remark 3
As an alternative to sharing a codebook of potentially infinite size, it is also possible for the BS and the users to utilize common randomness to generate random codewords as they are needed. In a sense, the codewords can be thought of as the output of a random hash function from the user indices $[n]$ to the source alphabet $\mathcal{X}$ (according to the distribution (23)). In this paper, we assume either the availability of common randomness or the ability to share an infinite codebook between the BS and the users.
Remark 4
Throughout this paper, we assume that there are no errors in the activity detection phase, i.e., the BS knows the identities of the active users perfectly. If instead, some inactive users are incorrectly detected as active, the sources intended for them would simply not be received. On the other hand, if some active users are missed in the activity detection phase, the encoding scheme proposed in this paper would cause these users to decode to erroneous source values. To remedy this latter more serious scenario, the BS can positively and negatively acknowledge the detected active and inactive users (at a cost of additional downlink bits.) The design of codes for acknowledgment has been studied in [16].
II-C Achievable Common Message Length
In this section, we provide upper bounds on the optimal common message length $R^{*}$ of the downlink message for massive random access by analyzing the entropy of the encoder output $H(g_{\mathbf{m}^{*}}(\mathbf{X},\mathbf{A}))$, when
  • Sources ${\mathbf{X}}=(X_{1},\cdots,X_{k})$ are distributed according to an exchangeable distribution $p(x_{1},\cdots,x_{k})$;
  • Activity pattern ${\mathbf{A}}$ is uniform among all possible activity patterns and is independent of $\mathbf{X}$;
  • The codebook $\mathbf{m}$ is chosen from an ensemble of codebooks generated according to i.i.d. mixture distribution $q(\mathbf{x}|\theta)r(\theta)$.
@@ II-C Achievable Common Message Length
Consider a massive random access scenario with a total of $n$ users and a random subset of $k$ active users. Let sources $\mathbf{X}=(X_{1},\cdots,X_{k})$ take values in a finite or countable set $\mathcal{X}^{k}$ with exchangeable distribution $p(\mathbf{x})$. The minimum common message length $R^{*}$ is bounded from above as
$R^{*}<\min_{q\in\mathcal{Q}}H(\mathbf{X})+D(p\|q)+\log(H(\mathbf{X})+D(p\|q)+1)+2,\$ (25)
where $\mathcal{Q}$ is the family of all i.i.d. mixture distributions on $\mathcal{X}^{k}$ defined as in (23). Further, the minimum common message length $R^{*}$ is also bounded from above as:
$R^{*}<\min_{q\in\mathcal{Q}}H(\mathbf{X})+D(p\|q)+\log\left(\log\left(\frac{\bar{q}_{\text{max}}}{\bar{q}_{\text{min}}}\right)+2\right)+4,$ (26)
@@
$\bar{q}_{\text{max}}=\sup_{\mathbf{x}\in\mathcal{X}^{k}:p(\mathbf{x})>0}q(\mathbf{x});\quad\bar{q}_{\text{min}}=\inf_{\mathbf{x}\in\mathcal{X}^{k}:p(\mathbf{x})>0}q(\mathbf{x}).$ (27)
@@
It is worth noting that the $D(p\|q)$ term is the divergence between the joint distributions $p(\mathbf{x})$ and $q(\mathbf{x})$, i.e.,
$D(p\|q)=\sum_{\mathbf{x}\in\mathcal{X}^{k}}p(x_{1},\cdots,x_{k})\log\left(\frac{p(x_{1},\cdots,x_{k})}{q(x_{1},\cdots,x_{k})}\right).$ (28)
In the setting of this paper, the above divergence term does not necessarily tensorize, as neither the distribution of the sources nor the distribution of the codebook entries are necessarily i.i.d.. However, the operational interpretation of the $D(p\|q)$ term is the same as in classical source coding; it corresponds to the extra cost of using a codebook constructed according to $q({\mathbf{x}})$ to compress sources with distribution $p({\mathbf{x}})$.
@@
Under the setting of Theorem 1, if the sources $\mathbf{X}$ have an i.i.d. mixture distribution $p(\mathbf{x})$, then
$R^{*}<H(\mathbf{X})+\min\bigg{\{}\log(H(\mathbf{X})+1)+2,\\ \log\left(\log\left(\frac{\bar{p}_{\text{max}}}{\bar{p}_{\text{min}}}\right)+2\right)+4\bigg{\}},$ (29)
@@
$\bar{p}_{\text{max}}=\sup_{\mathbf{x}\in\mathcal{X}^{k}:p(\mathbf{x})>0}p(\mathbf{x});\quad\bar{p}_{\text{min}}=\inf_{\mathbf{x}\in\mathcal{X}^{k}:p(\mathbf{x})>0}p(\mathbf{x}).$ (30)
Further, if the sources $\mathbf{X}$ are i.i.d. distributed according to $p(x)$, we have
$R^{*}<kH(p)+\min\bigg{\{}\log(kH(p)+1)+2,\\ \left.\log\left(k\log\left(\frac{p_{\text{max}}}{p_{\text{min}}}\right)+2\right)+4\right\},$ (31)
where
$p_{\text{max}}=\sup_{{x}\in\mathcal{X}:p({x})>0}p({x});\quad p_{\text{min}}=\inf_{{x}\in\mathcal{X}:p({x})>0}p({x}).$ (32)
@@
Setting $q(\mathbf{x})=p(\mathbf{x})$ is a natural choice for i.i.d. or i.i.d. mixture sources, because it maximizes the probability of matching codewords with the sources. This result shows that for i.i.d. and i.i.d. mixture sources, the proposed code construction and the proposed encoding and decoding scheme achieves the joint entropy of the sources to within at most an $O(\log(k))$ overhead. The overhead can be further reduced if the distribution is close to uniform. In particular, for an i.i.d. uniform distribution with $p_{\text{max}}=p_{\text{min}}$, the entropy can be achieved to within $O(1)$ bits.
An intuitive argument of why for i.i.d. sources, a matching i.i.d. codebook can approach the joint entropy of the sources is the following. When the codebook distribution and the distribution of the sources $p(\mathbf{x})$ match, for every realization of the sources $\mathbf{x}$, a randomly generated codeword would match $\mathbf{x}$ with a probability $p(\mathbf{x})$. Thus, the index of the first matching codeword follows a geometric distribution with parameter $p(\mathbf{x})$. The entropy of such a geometric distribution is essentially $-\log(p(\mathbf{x}))$ plus a constant. Then, when averaged over all realizations of $\mathbf{x}$, we can achieve a common message length of $\mathbb{E}[-\log(p(\mathbf{X}))]$, which is just the entropy of the sources $H(\mathbf{X})$, plus a constant.
To make the above argument work for any codebook distribution $q(\mathbf{x})$, which may not be the same as the source distribution $p(\mathbf{x})$, we recognize that in general the distribution of the first matching index is a mixture of geometric distributions, for which the exact computation of entropy is complicated. The proof in Appendix A uses a bounding technique due to [18], averaged over randomly generated codebooks, to show the existence of one codebook whose output has the desired entropy, up to an additive logarithmic term.